Find the x-intercept and y-intercept of 7x+2y=13
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Find the x-intercept and y-intercept of 7x+2y=13
----
x-intercept ?
Let y = 0; then x = 13/7
============================
y-intercept ?
Let x = 0; then y = 13/2
-----------------------------
Cheers,
Stan H.

Answer by solver91311(24713)



Substitute 0 for then calculate the value of that results. Say that value is , then the -intercept is the point

Substitute 0 for then calculate the value of that results. Say that value is , then the -intercept is the point

The values and will be ugly mixed number fractions in this problem.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism